OpenStudy (anonymous):

Solve \[\log(5x+14)=0\]

OpenStudy (lgbasallote):

do you have an idea what the base of log(5x+ 14) is?

OpenStudy (anonymous):

log(5x+14)=0 log(5x+14)=log_10^1 5x+14=1 5x=-13 x=-13/5

OpenStudy (anonymous):

no..not reallyy :/

OpenStudy (lgbasallote):

whenever a base is not given it is automatically assumed that the base is 10

OpenStudy (anonymous):

how can you turn 0 into log1??

OpenStudy (anonymous):

0 = log_(10)1

OpenStudy (lgbasallote):

so if you turn that into exponential function you'll get 10^0 = 5x + 14 anything raised to 0 is 1 so it becomes 1 = 5x + 14

OpenStudy (anonymous):

ok, i see now, thanks people :D